# Time Makes Space: Emergence of Place Fields in Networks Encoding Temporally Continuous Sensory Experiences

### v2

@@ 1 Introduction
The hippocampus is a brain region that plays a critical role in both spatial navigation knierimHippocampus2015 ; moserPlaceCellsGrid2015 ; moserSpatialRepresentationHippocampal2017 and episodic memory squireHippocampusMemorySpace1991 ; corkinWhatNewAmnesic2002 ; squireLegacyPatientNeuroscience2009 ; hasselmoHowWeRemember2011 ; hasselmoThetaRhythmEncoding2014 . Researchers have therefore sought to understand the neural substrates of these forms of memory in the hippocampus, as well as the extent to which these roles are interrelated eichenbaumCanWeReconcile2014 ; hasselmoModelsSpatialTemporal2017 ; eichenbaumIntegrationSpaceTime2017 ; whittingtonTolmanEichenbaumMachineUnifying2019 . A key role is played by place cells – so named because they fire in specific spatial locations within an environment okeefeHippocampusSpatialMap1971 . Place cells have spatially constrained firing fields, and “remap” firing patterns in response to significant contextual changes (especially in hippocampal subregion CA3) mullerEffectsChangesEnvironment1987 ; bostockExperienceDependentModifications1991 ; jezekThetapacedFlickeringPlacecell2011 ; solstadPlaceCellRate2014 ; leutgebRemappingDiscriminateContexts2014 ; markThetaPacedFlickering2017 , including motion to a new environment okeefeHippocampusCognitiveMap1978 and modified behavioral context or sensory cues almePlaceCellsHippocampus2014 ; kubieHippocampalRemappingPhysiological2020 . After remapping, they form almost orthogonal representations almePlaceCellsHippocampus2014 ; posaniFunctionalConnectivityModels2017 . These findings suggest that CA3 place cell network may play a key role in maintaining both positional and contextual information markThetaPacedFlickering2017 ; posaniIntegrationMultiplexingPositional2018 , while also supporting a large capacity for memory storage battistaCapacityResolution2020 . Meanwhile, others have suggested that hippocampal CA3 is a pattern completion and separation device marrSimpleMemoryTheory1971 ; yassaPatternSeparationHippocampus2011 ; rollsMechanismsPatternCompletion2013 ; knierimTrackingFlowHippocampal2016 that supports the storage and retrieval of episodic memories from partial cues rollsComputationalTheoryEpisodic2010 ; hasselmoHowWeRemember2011 ; knierimTrackingFlowHippocampal2016 . This claim is supported by the extensive recurrent collaterals in CA3 liHippocampalCA3Network1994 ; witterIntrinsicExtrinsicWiring2007 ; hasselmoHowWeRemember2011 , which could allow it to act as an autoassociative network retrieving a complete memory fron partial cues mcnaughtonHippocampalSynapticEnhancement1987 ; hasselmoEncodingRetrievalEpisodic1998 ; hasselmoHowWeRemember2011 ; knierimTrackingFlowHippocampal2016 .
@@ 1.1 Simulated trials and the artificial agent
An artificial agent explores simulated room(s) defined by fixed WSM signals, as defined in Section 2.2. The agent updates its behavior and samples sensory experience every 50ms. To realistically model rodent movement trajectories, we simulate a random walk, where artificial agents explore the room in a trajectory with random small drifts representing an animal’s exploration interest: at every timestep, the agent randomly changes its movement speed and direction with a small probability, 0.2 and 0.3, respectively. The change in speed is randomly chosen from an $\mathcal{N}(5,1)$ cm/s normal distribution, and the drift direction is randomly chosen from $\mathcal{N}(0,0.05)$ rad/s.
@@ 1.2 Training Recurrent Autoencoder (RAE)
The network is trained with mean-squared error without strict sparseness constraints. The loss function is:
$\mathcal{L}=\frac{\lambda_{mse}}{D\cdot T\cdot B}\sum_{d,t,b}^{D,T,B}(\hat{e}_{d,t,b}-e_{d,t,b})^{2}+\frac{\lambda_{fr}}{N}\sum_{n}^{N}\left(\sum_{t,b}^{T,B}r_{n,t,b}\right)^{2}$ (5)
$D$ is the dimension of the Experience Vector (EV), $T$ measures timesteps within a memory segment, $B$ indicates the number of batches, and $N$ counts nodes in the hidden layer. The term $\hat{e}_{d,t,b}$ is the $d$-th entry of the reconstructed EV at timestep $t$ in batch $b$, while $e_{d,t,b}$ is the equivalent entry of the ground truth. The second component of the loss, $r_{n,t,b}$, denotes the firing rate of the $n$-th neuron at time $t$ within batch $b$. Unless otherwise stated, we set $\lambda_{mse}=0$ and $\lambda_{fr}=200$ for our experiments.

### v3

@@ 1 Introduction
The hippocampus is a brain region that plays a critical role in both spatial navigation knierimHippocampus2015 ; moserPlaceCellsGrid2015 ; moserSpatialRepresentationHippocampal2017 and episodic memory squireHippocampusMemorySpace1991 ; corkinWhatNewAmnesic2002 ; squireLegacyPatientNeuroscience2009 ; hasselmoHowWeRemember2011 ; hasselmoThetaRhythmEncoding2014 . Researchers have therefore sought to understand the neural substrates of these forms of memory in the hippocampus, as well as the extent to which these roles are interrelated eichenbaumCanWeReconcile2014 ; hasselmoModelsSpatialTemporal2017 ; eichenbaumIntegrationSpaceTime2017 ; whittingtonTolmanEichenbaumMachineUnifying2019 . A key role is played by place cells – so named because they fire in specific spatial locations within an environment okeefeHippocampusSpatialMap1971 . Place cells have spatially constrained firing fields, and “remap” firing patterns in response to significant contextual changes (especially in hippocampal subregion CA3) mullerEffectsChangesEnvironment1987 ; bostockExperienceDependentModifications1991 ; jezekThetapacedFlickeringPlacecell2011 ; solstadPlaceCellRate2014 ; leutgebRemappingDiscriminateContexts2014 ; markThetaPacedFlickering2017 , including motion to a new environment okeefeHippocampusCognitiveMap1978 and modified behavioral context or sensory cues almePlaceCellsHippocampus2014 ; kubieHippocampalRemappingPhysiological2020 . After remapping, they form almost orthogonal representations almePlaceCellsHippocampus2014 ; posaniFunctionalConnectivityModels2017 . These findings suggest that CA3 place cell network may play a key role in maintaining both positional and contextual information markThetaPacedFlickering2017 ; posaniIntegrationMultiplexingPositional2018 , while also supporting a large capacity for memory storage PhysRevLett.124.048302 . Meanwhile, others have suggested that hippocampal CA3 is a pattern completion and separation device marrSimpleMemoryTheory1971 ; yassaPatternSeparationHippocampus2011 ; rollsMechanismsPatternCompletion2013 ; knierimTrackingFlowHippocampal2016 that supports the storage and retrieval of episodic memories from partial cues rollsComputationalTheoryEpisodic2010 ; hasselmoHowWeRemember2011 ; knierimTrackingFlowHippocampal2016 . This claim is supported by the extensive recurrent collaterals in CA3 liHippocampalCA3Network1994 ; witterIntrinsicExtrinsicWiring2007 ; hasselmoHowWeRemember2011 , which could allow it to act as an autoassociative network retrieving a complete memory fron partial cues mcnaughtonHippocampalSynapticEnhancement1987 ; hasselmoEncodingRetrievalEpisodic1998 ; hasselmoHowWeRemember2011 ; knierimTrackingFlowHippocampal2016 .
@@ 1.1 Simulated trials and the artificial agent
An artificial agent explores simulated room(s) defined by fixed WSM signals, as described in Section 2.2. The agent updates its location at each discretized timestep (50ms) to simulate rodent movement behavior. This simulated trajectory is then used to sample corresponding experience vectors from the simulated WSM maps.
To realistically model rodent movement trajectories, we simulate a biased random walk. At the beginning of the simulation, the agent samples an initial movement speed and direction, as well as a small directional drift. The directional drift simulates the effect of momentum when an animal turns and moves to a new location, causing its heading to remain partially influenced by its previous direction. At each timestep, the agent biases its movement direction by the current drift, then moves along that direction with the given speed. With small probabilities, 0.2 for speed and 0.3 for directional drift, the agent resamples its movement parameters. The new speed is drawn from a $\mathcal{N}(5,1)$ cm/s distribution, and the new drift from $\mathcal{N}(0,0.05)$ rad/s. Once updated, these parameters govern motion until they are resampled again.
@@ 1.2 Training Recurrent Autoencoder (RAE)
The network is trained with mean-squared error combined with a soft firing rate regularization, without enforcing strict sparseness constraints. The loss function is:
$\mathcal{L}=\frac{\lambda_{mse}}{D\cdot T\cdot B}\sum_{d,t,b}^{D,T,B}(\hat{e}_{d,t,b}-e_{d,t,b})^{2}+\frac{\lambda_{fr}}{N}\sum_{n}^{N}\left(\frac{1}{T\cdot B}\sum_{t,b}^{T,B}r_{n,t,b}\right)^{2}$ (5)
$D$ is the dimension of the Experience Vector (EV), $T$ measures timesteps within a memory segment, $B$ indicates the number of batches, and $N$ counts nodes in the hidden layer. The term $\hat{e}_{d,t,b}$ is the $d$-th entry of the reconstructed EV at timestep $t$ in batch $b$, while $e_{d,t,b}$ is the corresponding entry of the ground truth. The second component of the loss, $r_{n,t,b}$, denotes the firing rate of the $n$-th neuron at time $t$ within batch $b$. Unless otherwise stated, we set $\lambda_{mse}=1$ and $\lambda_{fr}=200$ for our experiments.